### v1

# Enhancing Depression-Diagnosis-Oriented Chat with Psychological State Tracking

Abstract
Depression-diagnosis-oriented chat aims to guide patients in self-expression to collect key symptoms for depression detection. Recent work focuses on combining task-oriented dialogue and chitchat to simulate the interview-based depression diagnosis. Whereas, these methods can not well capture the changing information, feelings, or symptoms of the patient during dialogues. Moreover, no explicit framework has been explored to guide the dialogue, which results in some useless communications that affect the experience. In this paper, we propose to integrate Psychological State Tracking (POST) within the large language model (LLM) to explicitly guide depression-diagnosis-oriented chat. Specifically, the state is adapted from a psychological theoretical model, which consists of four components, namely Stage, Information, Summary and Next. We fine-tune an LLM model to generate the dynamic psychological state, which is further used to assist response generation at each turn to simulate the psychiatrist. Experimental results on the existing benchmark show that our proposed method boosts the performance of all subtasks in depression-diagnosis-oriented chat.
1 Introduction
Depression remains an escalating mental health threat globally, due to the severe scarcity and limited access to professionals Sharma et al. (2020); Pérez et al. (2023). To alleviate such situations, computational methods are emerging tools for large-scale depression detection and diagnosis Zhang et al. (2021); Pacheco-Lorenzo et al. (2021). Conversations agents related to mental health become a promising technological solution for early depression detection unobtrusively, due to the traditional detection mechanisms being invasive Chaytor and Schmitter-Edgecombe (2003). The most popular implementation of chatbots is based on self-rating scales Jaiswal et al. (2019); Arrabales (2020) or diagnostic criteria Philip et al. (2017) by calibrating conversational interactions to cover the targeted topics on the scales, which in turn limits patients to express their feelings Yao et al. (2022). Therefore, there is still a significant gap for such systems to provide chat-like depression detection in clinical settings.
In practical psychotherapy, psychiatrists calibrate the dialogue flow to gather and synthesize key symptom information from patients while providing appropriate strategies like emotional support. To simulate this process, Yao et al. (2022) defined this type of dialogue as Task-oriented Chat and collected the first dialogue dataset D for depression diagnosis. However, existing work mostly focuses on shallow heuristic attempts such as predicting topics and generating empathetic responses, falling short of capturing the changing information, feelings or symptoms of the patient during dialogues.
Recently, Large Language Models (LLMs) have achieved remarkable success in various text reasoning tasks Kojima et al. (2022); Bang et al. (2023). In the field of psychology, ChatGPT and GPT-4 have shown promising performance in attributing mental states Bubeck et al. (2023); Chen et al. (2023c). While initial researchers envisioned further harnessing this capability for complex psychological tasks, the majority still focused on developing chatbots for emotional support purposes. Moreover, no explicit framework has been explored to guide the dialogue, which results in some useless communications that affect the experience.
To enhance depression-diagnosis-oriented chat, we propose the Psychological State Tracking (POST) to link the patient’s current symptom with the doctor’s next strategy. Inspired by the ABC Model Ellis (1991) in Cognitive-behavioral therapy (CBT), we define POST into four components: Stage, Information, Summary and Next. Figure 1 shows an illustrative example. First, we figure out at which stage of the current diagnosis procedure is; Then, we distinguish the key symptoms that the patient is exhibiting; After that, we document the current diagnostic results of the patient; Finally, in Next, we introduce an auxiliary targeted stimulus prompt to align with specific counseling strategies. We jointly optimize the POST model and the response generation model by an LLM. We conduct comprehensive experiments using the trained LLM. Experimental results on the existing benchmark show that our proposed method achieves the best performance of all subtasks in depression-diagnosis-oriented chat. Furthermore, psychological state tracking, as the explicit thought behind response generation, provides professional-compliant interpretability to the diagnostic process.
Overall, we make the following contributions:
We annotate a fine-grained dataset by augmenting the D4 dataset, which annotates the psychological state of each conversation round guided by Albert Ellis’ ABC Model.
We propose a joint model to explicitly guide depression-diagnosis-oriented chat, which integrates psychological state tracking into an LLM to learn the connection between patients’ state changes and doctors’ strategic planning.
Extensive experiments on the existing benchmark show that our proposed method boosts the performance of all subtasks in depression-diagnosis-oriented chat. In particular, psychological state tracking, as the explicit thought behind response generation, provides professional-compliant interpretability.
2 Related Work
2.1 Depression Diagnosis
Depression diagnosis aims to use diagnostic tools to identify symptoms and determine the severity of depression. Early research was conducted by psychiatrists in controlled settings through self-questionnaires, such as the PHQ-9 Kroenke et al. (2001) and GAD-7 Spitzer et al. (2006), to assess patients’ cognitive or emotional states. However, in face-to-face settings, individuals often hesitate to express their mental state. Some researchers explore using different network structures to automatically identify mental health status in social media content Bucur et al. (2023); Gui et al. (2019); An et al. (2020). The poor interactivity of these approaches also limits patients’ self-expression. Recent efforts focus on developing conversational agents to perform diagnosis in interactive environments Jaiswal et al. (2019); Arrabales (2020). Yao et al. (2022) propose to combine task-oriented dialogue and chitchat to simulate the interview-based depression diagnosis. Whereas, these methods struggle to well capture the changing information, feelings, or symptoms of the patient during the dialogue process. Moreover, there has been no exploration of an explicit framework to guide the response generation. Thus, we aim to explore a more personalized and professional depression diagnostic chatbot.
2.2 Large Language Models for Psychology
Large Language Models (LLMs) have demonstrated remarkable performance in interactions in zero-shot settings, igniting a fervent interest in their application within the field of psychology. Researchers adopt methodologies from human psychology to explore the behavior of LLMs Hagendorff (2023b), observing that the responses from LLMs often show a similarity to human responses in most scenarios Almeida et al. (2023). Additionally, LLMs have been utilized as auxiliary tools to advance specific psychological tasks Chen et al. (2023b); for example, Qin et al. (2023) utilized an LLM to develop an interpretable and interactive system for detecting depression. A few more recent works have begun to enhance the performance of LLMs with deeper professional psychological knowledge Chen et al. (2023c). These approaches not only enrich the capabilities of LLMs but also open new avenues for psychological research and applications. However, with the knowledge gap between the two communities, depending only on LLM-based systems for psychology can lead to unpredictable outcomes Hagendorff (2023a). Hence, we aim to further fine-tune LLMs combined the professional psychological theories to achieve more professional capabilities for psychology.
2.3 Dialogue State Tracking
Dialogue State Tracking (DST) is essential in task-oriented dialogue systems for monitoring conversation states. Previous studies have utilized pre-trained models to improve DST. For example, Lin et al. (2021) and Yu et al. (2022) employed large-scale pre-trained models for zero-shot cross-domain DST and for enhancing dialogue state tracking with external knowledge, respectively. Sun et al. (2022) revolutionizes dialogue state tracking with a Mentioned Slot Pool (MSP) to improve accuracy, while Xu et al. (2023) combines visit sequences and clinical notes through contrastive learning and adaptive language modeling to enhance diagnosis predictions. However, these approaches can not well handle complex dialogues and capture fine-grained semantic relationships.
The rise of large language models has led to advancements in DST methods, such as T5 Raffel et al. (2020) and GPT-3 Brown et al. (2020). Recent research focuses on parameter-efficient fine-tuning (PEFT) to enhance DST performance. Notable techniques include prompt tuning by Lee et al. (2021) and Yang et al. (2023), as well as prefix tuning Ma et al. (2023) and meta-learning Chen et al. (2023a). Despite these progresses, there remains a gap when compared to more advanced models like ChatGPT, and additional support from psychological theories is necessitated to accomplish our task. Therefore, our proposed methodology combines the strengths of open-source LLMs and fine-tuning techniques like LoRA Hu et al. (2021) to improve the fundamental generation ability, and utilizes the ABC model in Cognitive-behavioral therapy to design the DST framework.
3 Data Annotation
3.1 Annotation Procedure
We adapt the D Yao et al. (2022) dataset with additional psychological state annotations, which contains 1,339 clinically standardized conversations about depression. The clinical data can facilitate a generation and diagnosis process that closely simulates real-life clinical consultations for depression. However, the original data lacked tracking of the patient’s conditions. To transform the raw data into a sample set that can be used for psychological state tracking, we annotate the conversations following three steps: (1) First, the patient’s known information and current stage are captured in real-time as the conversation occurs. Special tokens representing the current stage and information are added to each utterance. (2) The annotation of summary and next action is then performed using in-house software. Our linguistic annotators are familiar with real-life clinical consultation scenarios, having completed training courses in depression diagnosis. We followed an annotation protocol defined by authoritative consultants for manual labeling. Subsequently, the manually labeled samples were used as input examples for GPT-4 to carry out the remaining annotations. (3) To ensure the annotation quality of GPT-4, we randomly sample 200 instances from the dataset annotated by GPT-4 and conduct manual verification. Out of these, 163 annotations are deemed accurate, resulting in an accuracy rate of 81.5%. The prompt template of annotation can be found in Appendix A.
3.2 Data Analysis
Source | Criteria | Total |
D | Dialogues | 1,339 |
Dialogue turns | 28,977 | |
Average turns per dialogue | 21.67 | |
Average tokens per dialogue | 577.12 | |
Average tokens per utterance | 13.31 | |
Average patient tokens per utterance | 11.87 | |
Average doctor tokens per utterance | 14.76 | |
POST | Stage-A per dialogue | 5.84 |
Stage-B per dialogue | 4.11 | |
Stage-C per dialogue | 11.72 | |
Average Info tokens per turn | 47.12 | |
Average Summary tokens per turn | 22.69 | |
Average POST tokens per turn | 75.81 |
Statistics
The basic statistics of the annotated dataset are shown in Table 1. The total Dialogues of the dataset is 1,339, while the total turns of dialogues is 28,977. Due to the data cleaning procedure, dialogue turns remaining may be less than the number of the original D dataset. The average doctor tokens per utterance is 14.76, which is approximately 3 tokens more than the average patient tokens, indicating that the doctors often speak more due to the need for consultation or empathetic consolation. The second part concerns the statistical analysis of annotated POSTs. The average number of POST tokens per turn is 75.81, suggesting that POSTs contain more information compared to utterances. Notably, the Info part, which contains patient information from dialogue history, is lengthier with an average length of 47.12. In contrast, the Summary, which involves further inference of the patient’s diagnostic results, is almost half the length of the Info part, averaging only 22.69.
Stage Analysis
The distribution of topics in different stages is illustrated in Figure 2. The chart reveals that Stage-A primarily focuses on Core and Behavior, aligning with its objective of identifying triggering events. In Stage-B, there is an emphasis on Empathetic Comfort and a notable increase in attention to Suicidal tendencies. This reflects the stage’s pivotal role in responding to patient disclosures and probing the beliefs of patients linked to identified symptoms and events. Conversely, Stage-C, as the terminal phase, primarily focuses on evaluating behavioral outcomes and intensifies screening to facilitate final diagnoses.
Figure 3 delineates the stage transitions in dialogues, highlighting the proportional changes every two turns starting from the second turn. Initially, dialogues commence with Stage-A, focusing on recent events potentially triggering the patient’s condition. After a thorough understanding of the precipitating events, the dialogue phase gradually transitions to Stage-B. This stage involves detailed inquiries into the patient’s thoughts and beliefs, and their attitudes towards these events, aiming to gain a deeper insight into the patient’s internal thoughts, including any tendencies towards depression or suicide. As the dialogue progresses, it shifts to Stage-C, where the clinician examines the patient’s core behaviors and symptoms, such as sleep and eating habits, to identify specific manifestations.
4 Method
Our approach formalizes depression-diagnosis-oriented chat by representing the user’s psychological state as a set of task attributes, as shown in Figure 1. The goal is to generate doctors’ probable responses based on the dialog context, taking into account the current state and next planning. As shown in Figure 4, following the task-specific fine-tuning paradigm, we build a joint model of psychological state tracking and response generation by equipping a transformer-based language backbone with functional modules.
4.1 Task Formulation
For a depression-diagnosis-oriented chat task, there is a turn dialogue between a patient and a doctor that can be represented as:
where is the patient’s psychological state at turn , is the patient’s utterance, and is the doctor’s response at turn . The psychological state of the patient represents the psychological diagnosis of thought, which consists of four subsequences . The entire depression-diagnosis-oriented chat procedure can be split into 4 subtasks: Psychological State Tracking predicts the patient’s current psychological state and the next action based on the dialogue context. Response Generation generates the most likely response based on the dialogue history and current psychological state. Similar to D, we jointly optimize the psychological state tracking model and the response generation model by an LLM. Dialogue Summary concentrates on generating summaries of symptoms, drawing from the complete dialogue context; Severity Classification assesses the severity of depression and suicide risk independently, utilizing information obtained from dialogue history and symptom summary.
4.2 Psychological State Tracking
Clinically, interview-based depression diagnosis need to collect and summarize key symptom information about one patient while providing a chat-like conversation experience. Psychiatrists need to obtain enough information from patients with multiple colloquial rhetoric based on the patient’s speech or written content. Thus, psychiatrists need to design the questioning logic between questions of symptoms from mild to severe during the consultation. To model such finer-grained relationships between patients’ symptoms and questions, we perform psychological state tracking, where the state consists of the following components.
Stage.
In clinical practice, a consultation follows a gradual in-depth manner and diagnosis topics consistently occur across turns. For example, doctors usually start by asking about core symptoms such as mood and interests, and then gradually turn to behavior symptoms. To perform a deep analysis of the topic transition, we first need to find out at which stage of the current diagnosis procedure. After an assessment, we summarize the dialogue stage into the context that served as the evidence base for asking about different types of symptoms.
Information.
This part aims to discover the psychological information that the patient is exhibiting in the dialogue history. Based on the presented symptoms, we derive the patient’s varying severity. For healthy individuals, the conversation typically manifests surface symptoms such as changes in sleep. As the condition worsens, patients tend to exhibit an increasing array of symptoms. By documenting the content and quantity of symptoms exhibited by the patient as a form of memory, we can gain a clearer understanding of the extent of their depression and devise strategies for planning.
Summary.
This subsequence aims to document the current diagnosis of the patient. Depression diagnoses are primarily employed for preliminary screening. For milder cases, empathetic strategies are generally used to encourage the patient’s self-expression, while for severe cases, immediate crisis intervention is required. Therefore, we have incorporated real-time diagnostic results into the process of depression diagnosis.
Next.
Given the inherent ambiguity in the depression diagnosis chat, defining clear planning as in other task-oriented conversations is challenging. To facilitate dialogue generation, we introduce an auxiliary targeted stimulus prompt to the psychological state. These prompts serve as nuanced, patient-specific hits and clues, guiding LLMs to align with specific counseling strategies.
We train LLMs to generate the above components in a psychological state. With psychological state tracking, we obtain a fully interpretable thought process for generating responses focused on depression diagnosis. As a diagnostic aid, such interpretability is crucial for professionals in clinical practice to substantiate the rationale behind the results, which can derive intricate thought patterns behind the entire diagnostic process from the generated thought processes.
4.3 Fine-tuning
In this work, we fine-tune ChatGLM3 111https://github.com/THUDM/ChatGLM3 with a parameter efficient approach, i.e., Low-Rank Adaptation (LoRA). LoRA maintains the weights of pre-trained LMs while introducing trainable rank decomposition matrices into each transformer layer, making it feasible to fine-tune LLMs with much fewer computational resources.
We fine-tune an LLM to track the psychological state and generate the response generation jointly, given the crafted example and the annotated label. Specifically, the objective is to predict the next token based on language modeling:
where represents the parameters for a language model and is the total turns of the dialogue. Ideally, the objective encourages the model to learn the target distribution by predicting tokens in the sequence. By placing the psychological state before the doctor’s response, the model learns to fuse the distribution from thought to response in an in-context language modeling manner. We only compute the loss of tokens on the psychological state and the doctor utterance .
5 Experiments
Model | BLEU-2 | ROUGE-L | METEOR | DIST-2 | Topic ACC. |
Transformer- | 7.28% | 0.21 | 0.1570 | 0.29 | - |
ChatGPT- | 10.01% | 0.19 | 0.3254 | 0.15 | - |
BART- | 19.29% | 0.35 | 0.2866 | 0.09 | - |
CPT- | 19.79% | 0.36 | 0.2969 | 0.07 | - |
Ours- | 30.85% | 0.47 | 0.4524 | 0.17 | - |
Transformer | 13.43% | 0.33 | 0.2620 | 0.04 | 36.82% |
ChatGPT | 17.52% | 0.30 | 0.4080 | 0.19 | 14.62% |
BART | 28.62% | 0.48 | 0.4053 | 0.07 | 59.56% |
CPT | 29.40% | 0.48 | 0.4142 | 0.06 | 59.77% |
Ours | 39.76% | 0.50 | 0.5305 | 0.11 | 56.98% |
Transformer* | 25.37% | 0.41 | 0.3905 | 0.04 | - |
ChatGPT* | 29.46% | 0.42 | 0.5158 | 0.23 | - |
BART* | 37.02% | 0.54 | 0.4920 | 0.07 | - |
CPT* | 37.45% | 0.54 | 0.4943 | 0.06 | - |
Ours* | 45.28% | 0.56 | 0.5794 | 0.10 | - |
5.1 Baselines
We leverage the original Transformer architecture Vaswani et al. (2017) fine-tuned on MedDialog dataset Zeng et al. (2020) as a foundational model. BART model Lewis et al. (2019) and the CPT model Shao et al. (2021) are employed as our primary baselines, with the same configuration as previous studies Yao et al. (2022). ChatGPT (gpt-3.5-turbo) 222https://openai.com/blog/chatgpt is listed as a baseline as well, which is optimized for chat by introducing Reinforcement Learning with Human Feedback (RLHF).
5.2 Implementation Details
Our model is implemented in PyTorch, using ChatGLM3-6b as the base architecture. In the supervised fine-tuning process, we apply LoRA to all linear layers of the model, where LoRA rank is set to 64. We set the batch size, max context length, and learning rate to 32, 1024, and 2e-4 respectively. The model is trained on one A800 GPU for 5 epochs, which costs about 8 hours.
5.3 Automatic Evaluation
Response Generation
Metrics like BLEU-2 Papineni et al. (2002), Rouge-L Lin (2004) and METEOR Banerjee and Lavie (2005) are employed to assess the response generation quality. In addition, we calculate DIST-2 Li et al. (2015) to demonstrate the diversity of responses. Tokenization is performed using the jieba333https://github.com/fxsjy/jieba library, following the same approach as in D.
Table 2 shows the results of the automatic evaluation for the response generation task. We have the following observations. First, without incorporating POSTs, the model achieves considerably better performances on BLEU-2, ROUGE-L and METEOR. It is mainly attributed to the larger scale of the language model and optimization for Chinese dialogue generation. Second, the integration of POSTs markedly enhanced diagnostic response capabilities, particularly in BLEU-2 and METEOR, by effectively assessing the diagnosis stage, summarizing patient information, and inferring potential diagnoses. By tracking psychological states, the POST guides the model in generating more rational and interpretable responses, and a further enhancement is observed while the golden POSTs are given. Third, The results from ChatGPT indicate relatively lower BLEU-2 and ROUGE-L scores due to the lack of dataset-specific fine-tuning. The introduction of POSTs notably improved ChatGPT’s performance by 7%, with substantial improvements upon using golden POSTs, underscoring the effectiveness of the POST strategy.
Summarization
Table 3 presents the results for dialog summary task. We provide the model with all the dialogue context and concatenate it with the patient’s portrait information. The model then generates a summary of the patient’s symptoms. The results show that our model surpasses both BART and CPT in text generation and also exhibits better diversity according to the DIST-2 scores.
Model | BLEU-2 | ROUGE-L | METEOR | DIST-2 |
BART | 16.44% | 0.26 | 0.25 | 0.19 |
CPT | 16.45% | 0.26 | 0.24 | 0.21 |
Ours | 18.32% | 0.30 | 0.29 | 0.26 |
Severity Classification
Results for depression severity and suicide risk severity classification are listed in Tables 4 and 5. To be consistent with D, we use average weighted precision, recall, and F1 scores computed by sklearn444https://scikit-learn.org to evaluate classification results.
The results demonstrate that our approach outperforms the baseline models in most scenarios, with a notable performance enhancement in 4-classification tasks when utilizing summaries compared to dialog history. However, the difference is less pronounced in the 2-classification tasks due to the increased complexity of the 4-classification tasks. Summaries provide concise and effective diagnostic information about the patient’s condition, which leads to a more substantial improvement in performance. Additionally, the variance in the 4-classification results based on dialog history is relatively high, which might be partly due to uneven label distribution in the dataset. Besides, the 2-classification accuracy of suicide severity based on dialog history have reached 90%, indicating this approach’s practical viability for diagnosing suicidal tendencies from dialog history.
Task | Input | Model | Precision | Recall | F1 |
2-class | dialog | BERT | 0.81 | 0.80 | 0.80 |
BART | 0.80 | 0.79 | 0.79 | ||
CPT | 0.79 | 0.78 | 0.78 | ||
Ours | 0.88 | 0.86 | 0.84 | ||
summary | BERT | 0.90 | 0.90 | 0.90 | |
BART | 0.89 | 0.89 | 0.89 | ||
CPT | 0.92 | 0.92 | 0.92 | ||
Ours | 0.91 | 0.91 | 0.91 | ||
4-class | dialog | BERT | 0.49 | 0.45 | 0.45 |
BART | 0.53 | 0.53 | 0.52 | ||
CPT | 0.49 | 0.47 | 0.46 | ||
Ours | 0.57 | 0.56 | 0.53 | ||
summary | BERT | 0.67 | 0.66 | 0.66 | |
BART | 0.68 | 0.67 | 0.66 | ||
CPT | 0.73 | 0.72 | 0.72 | ||
Ours | 0.79 | 0.76 | 0.76 |
Task | Input | Model | Precision | Recall | F1 |
2-class | dialog | BERT | 0.81 | 0.78 | 0.79 |
BART | 0.77 | 0.75 | 0.75 | ||
CPT | 0.84 | 0.82 | 0.82 | ||
Ours | 0.91 | 0.90 | 0.91 | ||
4-class | dialog | BERT | 0.72 | 0.64 | 0.66 |
BART | 0.70 | 0.66 | 0.65 | ||
CPT | 0.76 | 0.68 | 0.70 | ||
Ours | 0.79 | 0.71 | 0.73 |
5.4 Human Evaluation
To simulate realistic depression diagnosis scenarios for evaluation, we prompt ChatGPT to act as patients, based on patient backgrounds from the dataset. Correspondingly, Our model and baselines played the role of doctors, conducting diagnosis conversations with the patients. Each conversation consists of a minimum of 15 turns and will terminate at an appropriate round. Then, we assigned annotators with dialogue pairs to evaluate the performance of the doctor model following four aspects: 1) Fluency (Flu.) assesses the smoothness of the whole conversation; 2) Comforting (Com.) measures the ability to empathize and comfort; 3) Doctor-likeness (Doc.) gauges the adaptability in shifting topics based on the patient’s situation; 4) Engagingness (Eng.) measures if the model sustains attention throughout the conversation.
As shown in Table 6, our model surpasses the baseline in the majority of aspects, particularly in doctor-likeness and engagingness, indicating significant improvements compared to other models. This suggests that incorporating the POST enhances our model’s ability to emulate doctors’ interactions, capturing the patient’s state in real-time, and allowing for flexible topic transitions throughout the dialogue. However, we also note that our model did not outperform ChatGPT in terms of comforting. This could be attributed to ChatGPT’s ability to generate extensive comforting and advisory responses without specific tuning. The dialogue examples of these models can be found in Appendix B.
Comparisons | Aspect | Win | Lose | Tie |
Ours vs. CPT | Flu. | 44.7 | 21.2 | 34.1 |
Com. | 68.9 | 15.2 | 15.9 | |
Doc. | 62.1 | 14.4 | 23.5 | |
Eng. | 56.8 | 11.4 | 31.8 | |
Ours vs. ChatGPT | Flu. | 39.4 | 34.8 | 25.8 |
Com. | 33.3 | 55.3 | 11.4 | |
Doc. | 53.8 | 34.1 | 12.1 | |
Eng. | 41.7 | 28.8 | 29.5 |
5.5 Ablation Studies
To verify the effectiveness of our method, we conducted ablation studies. We used the model incorporating the golden POST as the baseline. Then, we removed each component of the POST in turn to study the effect of each part of the POST. The obtained results are demonstrated in Table 7.
Model | BLEU-2 | ROUGE-L | METEOR | DIST-2 |
Ours* | 45.28% | 0.56 | 0.5794 | 0.10 |
w/o Info | 44.43% | 0.54 | 0.5708 | 0.16 |
w/o Stage | 43.57% | 0.53 | 0.5681 | 0.16 |
w/o Sum | 44.25% | 0.54 | 0.5722 | 0.16 |
w/o Next | 36.24% | 0.45 | 0.5107 | 0.18 |
The study reveals that omitting the next topic (w/o Next) significantly impacts response generation, as it is directly correlated with the subsequent response’s relevance and effectiveness. This omission led to an increase in the DIST-2 value, suggesting a reduction in response content diversity due to the focused nature of topic-specific responses. Furthermore, excluding the stage (w/o Stage) notably affects the quality of generation, indicating the importance of the diagnostic dialogue phase in tailoring responses to be stage-appropriate and rational. In contrast, the absence of information (w/o Info) and summary (w/o Sum) components showed lesser impact on response generation quality, due to their roles in summarizing dialogue history rather than directly influencing responses. Nonetheless, these elements are crucial for their interpretability and utility in clinical settings, aiding doctors in summarizing patient symptoms and providing preliminary diagnostic insights, thereby facilitating further diagnosis and having a substantial role in clinical psychological consultations.
6 Conclusions
In this paper, we incorporate Psychological State Tracking (POST) within LLM to guide response generation for doctors during depression diagnosis consultations. In particular, the state is defined based on the Albert Ellis’ ABC Model in psychology, which illuminates a profound connection between patients’ information changes and doctors’ strategic planning. Extensive experiments show that the integration of psychological state tracking significantly enhances the performance of LLMs to generate responses in depression-diagnosis-oriented chat. Furthermore, our approach also provides explicit interpretations for using appropriate strategies in different situations to collect information or comfort patients for depression diagnosis. In the future, we will explore more specialized and fine-grained state tracking methods and incorporate patient personalized information to guide diagnosis-oriented chat.
Limitations
The main limitation of our approach lies in the evaluation metrics. In the field of diagnostic chat, traditional automatic evaluation methods may not accurately reflect the true effectiveness of the models. As highlighted by Liu et al. (2016), there exists a certain discrepancy between the results and human evaluations. Developing more comprehensive, considerate, and task-specific evaluation metrics remains a challenge to be addressed in future research.
Ethics Statement
In this study, we annotate the psychological state collaboratively by professional labors and GPT-4, followed by a thorough data cleaning and filtering process. This step ensured the removal of any private or sensitive information from the dataset.
Acknowledgements
We acknowledge the contributions of the SJTU X-LANCE Lab555https://x-lance.sjtu.edu.cn/en in constructing the original depression-diagnosis-oriented chat dataset (i.e., D). We cite the work and conduct further research by adapting the dataset and exploring LLM-based method from the psychological views.
Appendix A Prompt Template
Description: I want you to serve as a psychological counselor, and use rational emotive behavior therapy and ABC theory to mark the current patient situation and the next strategy. Please mark the input dialogues according to the requirements and examples in Chinese. The sequence numbers of the input dialogues must strictly correspond one-to-one with those of the outputs, with no merging or omissions allowed. Requirements: Definitions and marking requirements of four parts of POST: <Stage> The current diagnostic stage (Activating Event (A) / Belief (B) / Consequence (C), please enter the complete stage, and answers like "same as above" are not allowed). <Info> Current patient information (keep information from historical conversations as briefly as possible). <Sum> Diagnosis result of the patient’s condition (you need to refine it to the underlying disease). <Next> Next step strategy (choose the most suitable option from 11 strategies). Examples: Examples pre-labeled by humans Input: Input data Output:
Appendix B Case Study
Figures 5, 6, 7 respectively display dialogue examples of our model, CPT, and ChatGPT under the same patient background, where these models play the role of doctors. In the diagram, statements of empathetic consolation are highlighted with green wavy lines, while inquiries of diagnosis are underlined in blue. It is evident from the illustration that, in comparison to other models, our model offers more natural empathetic consolation and topics of inquiry that are closely aligned with the diagnostic state. The responses from CPT tend to be relatively rigid, with less empathetic consolation and more straightforward, singular diagnostic inquiries. Although ChatGPT provides commendable empathetic consolation, its replies often are too lengthy, containing extensive analysis and advice, yet they do not closely correlate with the diagnostic content, lacking the distinctive characteristics of a doctor.

### v2

@@
 Recent work focuses on combining task-oriented dialogue and chitchat to simulate the interview-based depression diagnosis.
-Whereas, these methods can not well capture the changing information, feelings, or symptoms of the patient during dialogues.
-Moreover, no explicit framework has been explored to guide the dialogue, which results in some useless communications that affect the experience.
+However, these methods can not well capture the changing information, feelings, or symptoms of the patient during dialogues.
+Moreover, no explicit framework has been explored to guide the dialogue, resulting in some ineffective communications that impact the experience.
 In this paper, we propose to integrate Psychological State Tracking (POST) within the large language model (LLM) to explicitly guide depression-diagnosis-oriented chat.
-Specifically, the state is adapted from a psychological theoretical model, which consists of four components, namely Stage, Information, Summary and Next.
+Specifically, the state is adapted from a psychological theoretical model, which consists of four components: Stage, Information, Summary, and Next.
 We fine-tune an LLM model to generate the dynamic psychological state, which is further used to assist response generation at each turn to simulate the psychiatrist.
 Experimental results on the existing benchmark show that our proposed method boosts the performance of all subtasks in depression-diagnosis-oriented chat.
-1 Introduction
-Depression remains an escalating mental health threat globally, due to the severe scarcity and limited access to professionals Sharma et al.
-(2020); Pérez et al.
-(2023).
-To alleviate such situations, computational methods are emerging tools for large-scale depression detection and diagnosis Zhang et al.
-(2021); Pacheco-Lorenzo et al.
-(2021).
-Conversations agents related to mental health become a promising technological solution for early depression detection unobtrusively, due to the traditional detection mechanisms being invasive Chaytor and Schmitter-Edgecombe (2003).
-The most popular implementation of chatbots is based on self-rating scales Jaiswal et al.
-(2019); Arrabales (2020) or diagnostic criteria Philip et al.
-(2017) by calibrating conversational interactions to cover the targeted topics on the scales, which in turn limits patients to express their feelings Yao et al.
-(2022).
-Therefore, there is still a significant gap for such systems to provide chat-like depression detection in clinical settings.
-In practical psychotherapy, psychiatrists calibrate the dialogue flow to gather and synthesize key symptom information from patients while providing appropriate strategies like emotional support.
+Keywords:
+Depression diagnosis chat Dialogue state tracking Large language models Dialogue systems Psychology.1 Introduction
+Depression remains an escalating mental health threat globally, due to the severe scarcity and limited access to professionals.
+To alleviate such situations, conversational agents become a promising solution for early depression detection, due to the traditional detection mechanisms being invasive [4].
+In practical psychotherapy, psychiatrists dynamically adjust the dialogue flow to effectively collect symptoms from patients, while providing appropriate intervention strategies such as emotional support.
 To simulate this process, Yao et al.
-(2022) defined this type of dialogue as Task-oriented Chat and collected the first dialogue dataset D for depression diagnosis.
+[25] defined this kind of dialogue as Task-oriented Chat and collected the first dialogue dataset D4 for depression diagnosis.
 However, existing work mostly focuses on shallow heuristic attempts such as predicting topics and generating empathetic responses, falling short of capturing the changing information, feelings or symptoms of the patient during dialogues.
-Recently, Large Language Models (LLMs) have achieved remarkable success in various text reasoning tasks Kojima et al.
-(2022); Bang et al.
-(2023).
-In the field of psychology, ChatGPT and GPT-4 have shown promising performance in attributing mental states Bubeck et al.
-(2023); Chen et al.
-(2023c).
-While initial researchers envisioned further harnessing this capability for complex psychological tasks, the majority still focused on developing chatbots for emotional support purposes.
-Moreover, no explicit framework has been explored to guide the dialogue, which results in some useless communications that affect the experience.
+Recently, Large Language Models (LLMs) have achieved remarkable success in various text reasoning tasks.
+In the field of psychology, ChatGPT [17] and GPT-4 [15] have shown promising performance in attributing mental states [2].
+While researchers have envisioned further harnessing this capability for complex psychological tasks, the majority still focus on developing chatbots for emotional support purposes.
+Moreover, no explicit framework has been explored to guide the dialogue, creating a gap between LLMs and psychiatrists, especially regarding their lack of skill in questioning [6].
 To enhance depression-diagnosis-oriented chat, we propose the Psychological State Tracking (POST) to link the patient’s current symptom with the doctor’s next strategy.
-Inspired by the ABC Model Ellis (1991) in Cognitive-behavioral therapy (CBT), we define POST into four components: Stage, Information, Summary and Next.
-Figure 1 shows an illustrative example.
-First, we figure out at which stage of the current diagnosis procedure is; Then, we distinguish the key symptoms that the patient is exhibiting; After that, we document the current diagnostic results of the patient; Finally, in Next, we introduce an auxiliary targeted stimulus prompt to align with specific counseling strategies.
+Inspired by Albert Ellis’ ABC Model [8] in Cognitive-behavioral therapy (CBT), we define the psychological state with four components: Stage, Information, Summary and Next.
+First, we figure out at which stage of the current diagnosis procedure is; Then, we distinguish the key symptoms information that the patient is exhibiting; After that, we document the current diagnostic summary of the patient; Finally, in Next, we introduce a targeted prompt to align with specific counseling strategies.
 We jointly optimize the POST model and the response generation model by an LLM.
-We conduct comprehensive experiments using the trained LLM.
 Experimental results on the existing benchmark show that our proposed method achieves the best performance of all subtasks in depression-diagnosis-oriented chat.
 Furthermore, psychological state tracking, as the explicit thought behind response generation, provides professional-compliant interpretability to the diagnostic process.
-Overall, we make the following contributions:
-We annotate a fine-grained dataset by augmenting the D4 dataset, which annotates the psychological state of each conversation round guided by Albert Ellis’ ABC Model.
+The main contributions of our work are as follows:
+We annotate a fine-grained dataset by augmenting the D4 dataset, which annotates the psychological state of each conversation round guided by the ABC Model.
 We propose a joint model to explicitly guide depression-diagnosis-oriented chat, which integrates psychological state tracking into an LLM to learn the connection between patients’ state changes and doctors’ strategic planning.
-Extensive experiments on the existing benchmark show that our proposed method boosts the performance of all subtasks in depression-diagnosis-oriented chat.
-In particular, psychological state tracking, as the explicit thought behind response generation, provides professional-compliant interpretability.
+Extensive experiments on the existing benchmark show that our proposed method boosts the performance of depression-diagnosis-oriented chat.
+In particular, the psychological state tracking serves as an explicit thought to provide interpretability for response generation.
 2 Related Work
 2.1 Depression Diagnosis
-Depression diagnosis aims to use diagnostic tools to identify symptoms and determine the severity of depression.
-Early research was conducted by psychiatrists in controlled settings through self-questionnaires, such as the PHQ-9 Kroenke et al.
-(2001) and GAD-7 Spitzer et al.
-(2006), to assess patients’ cognitive or emotional states.
+Depression diagnosis aims to use diagnostic tools to identify symptoms and determine the severity of depression [7, 13].
+Early research was conducted by psychiatrists in controlled settings through self-questionnaires, such as the PHQ-9 [10] and GAD-7 [20], to assess patients’ cognitive or emotional states.
 However, in face-to-face settings, individuals often hesitate to express their mental state.
-Some researchers explore using different network structures to automatically identify mental health status in social media content Bucur et al.
-(2023); Gui et al.
-(2019); An et al.
-(2020).
+Some researchers explore using different network structures to automatically identify mental health status in social media content [3].
 The poor interactivity of these approaches also limits patients’ self-expression.
-Recent efforts focus on developing conversational agents to perform diagnosis in interactive environments Jaiswal et al.
-(2019); Arrabales (2020).
 Yao et al.
-(2022) propose to combine task-oriented dialogue and chitchat to simulate the interview-based depression diagnosis.
+[25] proposed to combine task-oriented dialogue and chitchat to simulate the interview-based depression diagnosis.
+Seo et al.
+[18] integrated depression diagnosis into emotional support conversation to improve diagnosis ability.
 Whereas, these methods struggle to well capture the changing information, feelings, or symptoms of the patient during the dialogue process.
@@
 Thus, we aim to explore a more personalized and professional depression diagnostic chatbot.
-2.2 Large Language Models for Psychology
-Large Language Models (LLMs) have demonstrated remarkable performance in interactions in zero-shot settings, igniting a fervent interest in their application within the field of psychology.
-Researchers adopt methodologies from human psychology to explore the behavior of LLMs Hagendorff (2023b), observing that the responses from LLMs often show a similarity to human responses in most scenarios Almeida et al.
-(2023).
-Additionally, LLMs have been utilized as auxiliary tools to advance specific psychological tasks Chen et al.
-(2023b); for example, Qin et al.
-(2023) utilized an LLM to develop an interpretable and interactive system for detecting depression.
-A few more recent works have begun to enhance the performance of LLMs with deeper professional psychological knowledge Chen et al.
-(2023c).
-These approaches not only enrich the capabilities of LLMs but also open new avenues for psychological research and applications.
-However, with the knowledge gap between the two communities, depending only on LLM-based systems for psychology can lead to unpredictable outcomes Hagendorff (2023a).
-Hence, we aim to further fine-tune LLMs combined the professional psychological theories to achieve more professional capabilities for psychology.
 2.2 Dialogue State Tracking
-Dialogue State Tracking (DST) is essential in task-oriented dialogue systems for monitoring conversation states.
+Dialogue State Tracking (DST) [22] is essential in task-oriented dialogue systems for monitoring conversation states.
 Previous studies have utilized pre-trained models to improve DST.
-For example, Lin et al.
-(2021) and Yu et al.
-(2022) employed large-scale pre-trained models for zero-shot cross-domain DST and for enhancing dialogue state tracking with external knowledge, respectively.
+For example, Wu et al.
+[23] employed large-scale pre-trained models for zero-shot DST and enhanced dialogue state tracking with intricate updating strategies.
 Sun et al.
-(2022) revolutionizes dialogue state tracking with a Mentioned Slot Pool (MSP) to improve accuracy, while Xu et al.
-(2023) combines visit sequences and clinical notes through contrastive learning and adaptive language modeling to enhance diagnosis predictions.
-However, these approaches can not well handle complex dialogues and capture fine-grained semantic relationships.
-The rise of large language models has led to advancements in DST methods, such as T5 Raffel et al.
-(2020) and GPT-3 Brown et al.
-(2020).
-Recent research focuses on parameter-efficient fine-tuning (PEFT) to enhance DST performance.
-Notable techniques include prompt tuning by Lee et al.
-(2021) and Yang et al.
-(2023), as well as prefix tuning Ma et al.
-(2023) and meta-learning Chen et al.
-(2023a).
-Despite these progresses, there remains a gap when compared to more advanced models like ChatGPT, and additional support from psychological theories is necessitated to accomplish our task.
-Therefore, our proposed methodology combines the strengths of open-source LLMs and fine-tuning techniques like LoRA Hu et al.
-(2021) to improve the fundamental generation ability, and utilizes the ABC model in Cognitive-behavioral therapy to design the DST framework.
+[21] revolutionize dialogue state tracking with a Mentioned Slot Pool (MSP) to improve accuracy.
+However, these approaches cannot effectively handle complex dialogues and capture fine-grained semantic relationships.
+The rise of large language models has led to advancements in DST methods.
+Recent research focuses on prompt learning [24], meta-learning [5], and LLM agents [14] to enhance DST performance.
+Despite these progressions, there remains a gap when compared to more advanced models like ChatGPT, and additional support from psychological theories is necessary to accomplish our task.
+Therefore, we designed the DST framework based on the ABC model of Cognitive-Behavioral Therapy and integrated it into large language models (LLMs) to guide the generation of responses for depression diagnosis.
+Fine-tuning techniques such as LoRA [9] are also applied to further improve the generation performance of the LLMs.
 3 Data Annotation
 3.1 Annotation Procedure
-We adapt the D Yao et al.
-(2022) dataset with additional psychological state annotations, which contains 1,339 clinically standardized conversations about depression.
+We adapt the D4 [25] dataset with additional psychological state annotations, which contains 1,339 clinically standardized conversations about depression.
 The clinical data can facilitate a generation and diagnosis process that closely simulates real-life clinical consultations for depression.
 However, the original data lacked tracking of the patient’s conditions.
-To transform the raw data into a sample set that can be used for psychological state tracking, we annotate the conversations following three steps: (1) First, the patient’s known information and current stage are captured in real-time as the conversation occurs.
-Special tokens representing the current stage and information are added to each utterance.
-(2) The annotation of summary and next action is then performed using in-house software.
-Our linguistic annotators are familiar with real-life clinical consultation scenarios, having completed training courses in depression diagnosis.
-We followed an annotation protocol defined by authoritative consultants for manual labeling.
-Subsequently, the manually labeled samples were used as input examples for GPT-4 to carry out the remaining annotations.
-(3) To ensure the annotation quality of GPT-4, we randomly sample 200 instances from the dataset annotated by GPT-4 and conduct manual verification.
-Out of these, 163 annotations are deemed accurate, resulting in an accuracy rate of 81.5%.
-The prompt template of annotation can be found in Appendix A.
+To transform the raw data into a sample dataset that can be used for psychological state tracking, we annotate the conversations following three steps, as shown in Figure 1:
+First, we constructed an annotation framework to capture psychological states in real-time during the conversations.
+Special tokens were added to the beginning of each utterance to indicate the Current Stage based on Albert Ellis’ ABC Model, Patient’s Information, Diagnosis Summaries, and Next Strategies.
+The current stage refers to the three key stages in ABC Model: Stage-A focuses on identifying Activating events that trigger emotional responses; Stage-B centers on understanding the patient’s Beliefs and thought patterns regarding these events; and Stage-C evaluates the emotional and behavioral Consequences resulting from these beliefs.
+Second, we implemented fine-grained data annotation using a system developed based on LabelStudio 111https://labelstud.io.
+Three professional psychologists with expertise in clinical depression consultations were invited to perform the initial manual annotations following standard clinical protocols.
+These carefully annotated samples then served as seeds for GPT-4 to extend the annotation process to the remaining dataset, balancing annotation quality with efficiency while maintaining professional standards throughout the process.
+Third, we conducted thorough data cleaning and validation on the annotated dataset.
+To ensure the quality and reliability of GPT-4 annotations, our psychologists manually reviewed and verified all labels generated by GPT-4.
 3.2 Data Analysis
-Source | Criteria | Total |
-D | Dialogues | 1,339 |
-Dialogue turns | 28,977 | |
-Average turns per dialogue | 21.67 | |
-Average tokens per dialogue | 577.12 | |
-Average tokens per utterance | 13.31 | |
-Average patient tokens per utterance | 11.87 | |
-Average doctor tokens per utterance | 14.76 | |
-POST | Stage-A per dialogue | 5.84 |
-Stage-B per dialogue | 4.11 | |
-Stage-C per dialogue | 11.72 | |
-Average Info tokens per turn | 47.12 | |
-Average Summary tokens per turn | 22.69 | |
-Average POST tokens per turn | 75.81 |
-Statistics
-The basic statistics of the annotated dataset are shown in Table 1.
-The total Dialogues of the dataset is 1,339, while the total turns of dialogues is 28,977.
-Due to the data cleaning procedure, dialogue turns remaining may be less than the number of the original D dataset.
-The average doctor tokens per utterance is 14.76, which is approximately 3 tokens more than the average patient tokens, indicating that the doctors often speak more due to the need for consultation or empathetic consolation.
+[…]
+Dialogues | 1,339 | |
+Dialogue turns | 28,977 | ||
+Average turns per dialogue | 21.67 | ||
+Average tokens per dialogue | 577.12 | ||
+Average tokens per utterance | 13.31 | ||
+Average patient tokens per utterance | 11.87 | ||
+Average doctor tokens per utterance | 14.76 | ||
+POST | Stage-A per dialogue | 5.84 | |
+Stage-B per dialogue | 4.11 | ||
+Stage-C per dialogue | 11.72 | ||
+Average Info tokens per turn | 47.12 | ||
+Average Summary tokens per turn | 22.69 | ||
+Average POST tokens per turn | 75.81 | ||
+[…]
+3.2.1 Statistics
+The basic statistics of the annotated dataset are shown in Table 3.2.
+The total dialogues in the dataset are 1,339, while the total turns of dialogues are 28,977.
+Due to the data cleaning procedure, the remaining dialogue turns may be less than the number in the original D4 dataset.
+The average number of doctor tokens per utterance is 14.76, which is approximately 3 tokens more than the average number of patient tokens, indicating that doctors often speak more due to the need for consultation or empathetic consolation.
 The second part concerns the statistical analysis of annotated POSTs.
@@
 In contrast, the Summary, which involves further inference of the patient’s diagnostic results, is almost half the length of the Info part, averaging only 22.69.
-Stage Analysis
-The distribution of topics in different stages is illustrated in Figure 2.
+3.2.2 Stage Analysis
+The distribution of the next strategies in different stages is illustrated in Figure 2.
 The chart reveals that Stage-A primarily focuses on Core and Behavior, aligning with its objective of identifying triggering events.
-In Stage-B, there is an emphasis on Empathetic Comfort and a notable increase in attention to Suicidal tendencies.
-This reflects the stage’s pivotal role in responding to patient disclosures and probing the beliefs of patients linked to identified symptoms and events.
+In Stage-B, there is a notable increase in attention to Empathetic Comfort and Suicidal tendencies.
 Conversely, Stage-C, as the terminal phase, primarily focuses on evaluating behavioral outcomes and intensifies screening to facilitate final diagnoses.
-Figure 3 delineates the stage transitions in dialogues, highlighting the proportional changes every two turns starting from the second turn.
-Initially, dialogues commence with Stage-A, focusing on recent events potentially triggering the patient’s condition.
-After a thorough understanding of the precipitating events, the dialogue phase gradually transitions to Stage-B.
-This stage involves detailed inquiries into the patient’s thoughts and beliefs, and their attitudes towards these events, aiming to gain a deeper insight into the patient’s internal thoughts, including any tendencies towards depression or suicide.
-As the dialogue progresses, it shifts to Stage-C, where the clinician examines the patient’s core behaviors and symptoms, such as sleep and eating habits, to identify specific manifestations.
 4 Method
-Our approach formalizes depression-diagnosis-oriented chat by representing the user’s psychological state as a set of task attributes, as shown in Figure 1.
+Our approach formalizes depression-diagnosis-oriented chat by representing the user’s psychological state as a set of task attributes.
 The goal is to generate doctors’ probable responses based on the dialog context, taking into account the current state and next planning.
-As shown in Figure 4, following the task-specific fine-tuning paradigm, we build a joint model of psychological state tracking and response generation by equipping a transformer-based language backbone with functional modules.
+As shown in Figure 3, following the task-specific fine-tuning paradigm, we build a joint model for psychological state tracking and response generation by equipping a transformer-based language backbone with functional modules.
 4.1 Task Formulation
 For a depression-diagnosis-oriented chat task, there is a turn dialogue between a patient and a doctor that can be represented as:
-where is the patient’s psychological state at turn , is the patient’s utterance, and is the doctor’s response at turn .
-The psychological state of the patient represents the psychological diagnosis of thought, which consists of four subsequences .
-The entire depression-diagnosis-oriented chat procedure can be split into 4 subtasks: Psychological State Tracking predicts the patient’s current psychological state and the next action based on the dialogue context.
-Response Generation generates the most likely response based on the dialogue history and current psychological state.
-Similar to D, we jointly optimize the psychological state tracking model and the response generation model by an LLM.
-Dialogue Summary concentrates on generating summaries of symptoms, drawing from the complete dialogue context; Severity Classification assesses the severity of depression and suicide risk independently, utilizing information obtained from dialogue history and symptom summary.
+where is the patient’s utterance, and is the doctor’s response at turn .
+The entire depression-diagnosis-oriented chat procedure can be split into 2 subtasks: Psychological State Tracking predicts the patient’s current psychological state based on the dialogue context, which includes Stage, Information, Summary and Next.
+Response Generation generates the most likely response based on the dialogue history and current state.
+We jointly optimize the psychological state tracking model and the response generation model by an LLM.
 4.2 Psychological State Tracking
-Clinically, interview-based depression diagnosis need to collect and summarize key symptom information about one patient while providing a chat-like conversation experience.
-Psychiatrists need to obtain enough information from patients with multiple colloquial rhetoric based on the patient’s speech or written content.
-Thus, psychiatrists need to design the questioning logic between questions of symptoms from mild to severe during the consultation.
-To model such finer-grained relationships between patients’ symptoms and questions, we perform psychological state tracking, where the state consists of the following components.
-Stage.
-In clinical practice, a consultation follows a gradual in-depth manner and diagnosis topics consistently occur across turns.
+Clinically, interview-based depression diagnosis needs to collect and summarize key symptom information about the patient while providing a chat-like conversation experience.
+Psychiatrists need to design the questioning logic between questions of symptoms from mild to severe during the consultation.
+To model such fine-grained relationships between patients’ symptoms and questions, we perform psychological state tracking.
+The state consists of the following components:
+where stand for Stage, Information, Summary and Next of dialogue turn .
+Stage
+In clinical practice, a consultation follows a gradual in-depth manner and diagnosis strategies consistently occur across turns.
 For example, doctors usually start by asking about core symptoms such as mood and interests, and then gradually turn to behavior symptoms.
-To perform a deep analysis of the topic transition, we first need to find out at which stage of the current diagnosis procedure.
-After an assessment, we summarize the dialogue stage into the context that served as the evidence base for asking about different types of symptoms.
-Information.
-This part aims to discover the psychological information that the patient is exhibiting in the dialogue history.
-Based on the presented symptoms, we derive the patient’s varying severity.
+To perform a deep analysis of the strategy transition, we first need to find out at which stage of the current diagnosis procedure.
+After an assessment, we summarize the dialogue stage into Stage-A, Stage-B, or Stage-C, which represents Identifying Activating Events, Perceiving Patients’ Beliefs, and Assessing Consequences respectively.
+The process can be formulated as:
+where is the stage of previous turn, represents our POST model, denotes the parameters of the model, and represents the current dialogue history, which consists of turns of dialogue and patient’s utterance at turn :
+Information
+This part aims to discover the psychological information that exhibited in the dialogue history.
+Based on the presented symptoms, we derive the patient’s illness severity.
 For healthy individuals, the conversation typically manifests surface symptoms such as changes in sleep.
 As the condition worsens, patients tend to exhibit an increasing array of symptoms.
-By documenting the content and quantity of symptoms exhibited by the patient as a form of memory, we can gain a clearer understanding of the extent of their depression and devise strategies for planning.
-Summary.
-This subsequence aims to document the current diagnosis of the patient.
+Then we record the symptoms info by the following formulation:
+where represents the information of dialogue at turn .
+By documenting the symptoms exhibited by the patient as a form of memory, we can gain a clearer understanding of their depression status and design strategies for responses.
+Summary
+This subsequence aims to document the current diagnosis summary of the patient.
 Depression diagnoses are primarily employed for preliminary screening.
-For milder cases, empathetic strategies are generally used to encourage the patient’s self-expression, while for severe cases, immediate crisis intervention is required.
-Therefore, we have incorporated real-time diagnostic results into the process of depression diagnosis.
-Next.
-Given the inherent ambiguity in the depression diagnosis chat, defining clear planning as in other task-oriented conversations is challenging.
-To facilitate dialogue generation, we introduce an auxiliary targeted stimulus prompt to the psychological state.
-These prompts serve as nuanced, patient-specific hits and clues, guiding LLMs to align with specific counseling strategies.
-We train LLMs to generate the above components in a psychological state.
+For milder cases, empathetic strategies are generally used to encourage the patient’s self-expression.
+But for severe cases, immediate crisis intervention is required.
+Therefore, we have incorporated real-time diagnostic results into the process of depression diagnosis, formulated as:
+Next
+To facilitate dialogue generation, we introduce the Next strategy to guide the LLM, which considers the current stage, the patient’s symptom information, and the severity of depression.
+It determines which strategy and topic should be used in the next response to support further diagnosis:
+After the LLM finishes the generation for all psychological states, we prompt it with a combined pair of the states and the current dialogue history to generate the doctor’s response:
 With psychological state tracking, we obtain a fully interpretable thought process for generating responses focused on depression diagnosis.
-As a diagnostic aid, such interpretability is crucial for professionals in clinical practice to substantiate the rationale behind the results, which can derive intricate thought patterns behind the entire diagnostic process from the generated thought processes.
 4.3 Fine-tuning
-In this work, we fine-tune ChatGLM3 111https://github.com/THUDM/ChatGLM3 with a parameter efficient approach, i.e., Low-Rank Adaptation (LoRA).
+In this work, we fine-tune large language models with a parameter-efficient approach, i.e., Low-Rank Adaptation (LoRA).
 LoRA maintains the weights of pre-trained LMs while introducing trainable rank decomposition matrices into each transformer layer, making it feasible to fine-tune LLMs with much fewer computational resources.
-We fine-tune an LLM to track the psychological state and generate the response generation jointly, given the crafted example and the annotated label.
+We fine-tune an LLM to track the psychological state and generate the response jointly, given the crafted example and the annotated label.
 Specifically, the objective is to predict the next token based on language modeling:
@@
 5 Experiments
-Model | BLEU-2 | ROUGE-L | METEOR | DIST-2 | Topic ACC. |
-Transformer- | 7.28% | 0.21 | 0.1570 | 0.29 | - |
-ChatGPT- | 10.01% | 0.19 | 0.3254 | 0.15 | - |
-BART- | 19.29% | 0.35 | 0.2866 | 0.09 | - |
-CPT- | 19.79% | 0.36 | 0.2969 | 0.07 | - |
-Ours- | 30.85% | 0.47 | 0.4524 | 0.17 | - |
-Transformer | 13.43% | 0.33 | 0.2620 | 0.04 | 36.82% |
-ChatGPT | 17.52% | 0.30 | 0.4080 | 0.19 | 14.62% |
-BART | 28.62% | 0.48 | 0.4053 | 0.07 | 59.56% |
-CPT | 29.40% | 0.48 | 0.4142 | 0.06 | 59.77% |
-Ours | 39.76% | 0.50 | 0.5305 | 0.11 | 56.98% |
-Transformer* | 25.37% | 0.41 | 0.3905 | 0.04 | - |
-ChatGPT* | 29.46% | 0.42 | 0.5158 | 0.23 | - |
-BART* | 37.02% | 0.54 | 0.4920 | 0.07 | - |
-CPT* | 37.45% | 0.54 | 0.4943 | 0.06 | - |
-Ours* | 45.28% | 0.56 | 0.5794 | 0.10 | - |
-5.1 Baselines
-We leverage the original Transformer architecture Vaswani et al.
-(2017) fine-tuned on MedDialog dataset Zeng et al.
-(2020) as a foundational model.
-BART model Lewis et al.
-(2019) and the CPT model Shao et al.
-(2021) are employed as our primary baselines, with the same configuration as previous studies Yao et al.
-(2022).
-ChatGPT (gpt-3.5-turbo) 222https://openai.com/blog/chatgpt is listed as a baseline as well, which is optimized for chat by introducing Reinforcement Learning with Human Feedback (RLHF).
-5.2 Implementation Details
-Our model is implemented in PyTorch, using ChatGLM3-6b as the base architecture.
-In the supervised fine-tuning process, we apply LoRA to all linear layers of the model, where LoRA rank is set to 64.
-We set the batch size, max context length, and learning rate to 32, 1024, and 2e-4 respectively.
+5.1 Experimental Setups
+5.1.1 Baselines
+We leverage the CPT model [19] as our primary baseline, as it achieved the best performance in previous studies [25].
+And we use the same configuration as them.
+ChatGPT (gpt-3.5-turbo) 222https://openai.com/blog/chatgpt is also listed as a baseline, which performs well in most tasks.
+5.1.2 Implementation Details
+Our model uses ChatGLM3-6b333https://github.com/THUDM/ChatGLM3 as the base architecture, and is implemented in PyTorch.
+During the supervised fine-tuning process, we apply LoRA to all linear layers of the model, where LoRA rank is set to 64.
+We set the batch size, max context length, and learning rate to 32, 1024, and 2e-4, respectively.
 The model is trained on one A800 GPU for 5 epochs, which costs about 8 hours.
+[…]
+ChatGPT | / | 10.01% | 0.19 | 0.3254 | 0.15 | - |
+CPT | / | 19.79% | 0.36 | 0.2969 | 0.07 | - |
+ChatGLM3 | / | 30.85% | 0.47 | 0.4524 | 0.17 | - |
+ChatGPT | +POST | 17.52% | 0.30 | 0.4080 | 0.19 | 14.62% |
++7.51% | +0.11 | +0.0826 | +0.04 | - | ||
+CPT | +POST | 34.15% | 0.44 | 0.4609 | 0.06 | 57.12% |
++14.36% | +0.08 | +0.1640 | -0.01 | - | ||
+ChatGLM3 | +POST | 39.76% | 0.50 | 0.5305 | 0.11 | 56.98% |
++8.91% | +0.03 | +0.0781 | -0.06 | - | ||
+ChatGPT | +POST* | 29.46% | 0.42 | 0.5158 | 0.23 | - |
++19.45% | +0.23 | +0.1904 | +0.08 | - | ||
+CPT | +POST* | 41.94% | 0.55 | 0.5285 | 0.04 | - |
++22.15% | +0.19 | +0.2316 | -0.03 | - | ||
+ChatGLM3 | +POST* | 45.28% | 0.56 | 0.5794 | 0.10 | - |
++14.43% | +0.09 | +0.1270 | -0.07 | - | ||
+[…]
 5.2 Automatic Evaluation
-Response Generation
-Metrics like BLEU-2 Papineni et al.
-(2002), Rouge-L Lin (2004) and METEOR Banerjee and Lavie (2005) are employed to assess the response generation quality.
-In addition, we calculate DIST-2 Li et al.
-(2015) to demonstrate the diversity of responses.
-Tokenization is performed using the jieba333https://github.com/fxsjy/jieba library, following the same approach as in D.
+We evaluate the performance of turn-based response generation given the dialogue history in the dataset.
+Typical automatic metrics for text generation like BLEU-2 [16], Rouge-L [12] and METEOR [1] are employed to assess the response generation quality.
+In addition, we calculate DIST-2 [11] to demonstrate the diversity of generated responses.
 Table 2 shows the results of the automatic evaluation for the response generation task.
 We have the following observations.
-First, without incorporating POSTs, the model achieves considerably better performances on BLEU-2, ROUGE-L and METEOR.
-It is mainly attributed to the larger scale of the language model and optimization for Chinese dialogue generation.
-Second, the integration of POSTs markedly enhanced diagnostic response capabilities, particularly in BLEU-2 and METEOR, by effectively assessing the diagnosis stage, summarizing patient information, and inferring potential diagnoses.
-By tracking psychological states, the POST guides the model in generating more rational and interpretable responses, and a further enhancement is observed while the golden POSTs are given.
-Third, The results from ChatGPT indicate relatively lower BLEU-2 and ROUGE-L scores due to the lack of dataset-specific fine-tuning.
-The introduction of POSTs notably improved ChatGPT’s performance by 7%, with substantial improvements upon using golden POSTs, underscoring the effectiveness of the POST strategy.
-Summarization
-Table 3 presents the results for dialog summary task.
-We provide the model with all the dialogue context and concatenate it with the patient’s portrait information.
-The model then generates a summary of the patient’s symptoms.
-The results show that our model surpasses both BART and CPT in text generation and also exhibits better diversity according to the DIST-2 scores.
-Model | BLEU-2 | ROUGE-L | METEOR | DIST-2 |
-BART | 16.44% | 0.26 | 0.25 | 0.19 |
-CPT | 16.45% | 0.26 | 0.24 | 0.21 |
-Ours | 18.32% | 0.30 | 0.29 | 0.26 |
-Severity Classification
-Results for depression severity and suicide risk severity classification are listed in Tables 4 and 5.
-To be consistent with D, we use average weighted precision, recall, and F1 scores computed by sklearn444https://scikit-learn.org to evaluate classification results.
-The results demonstrate that our approach outperforms the baseline models in most scenarios, with a notable performance enhancement in 4-classification tasks when utilizing summaries compared to dialog history.
-However, the difference is less pronounced in the 2-classification tasks due to the increased complexity of the 4-classification tasks.
-Summaries provide concise and effective diagnostic information about the patient’s condition, which leads to a more substantial improvement in performance.
-Additionally, the variance in the 4-classification results based on dialog history is relatively high, which might be partly due to uneven label distribution in the dataset.
-Besides, the 2-classification accuracy of suicide severity based on dialog history have reached 90%, indicating this approach’s practical viability for diagnosing suicidal tendencies from dialog history.
-Task | Input | Model | Precision | Recall | F1 |
-2-class | dialog | BERT | 0.81 | 0.80 | 0.80 |
-BART | 0.80 | 0.79 | 0.79 | ||
-CPT | 0.79 | 0.78 | 0.78 | ||
-Ours | 0.88 | 0.86 | 0.84 | ||
-summary | BERT | 0.90 | 0.90 | 0.90 | |
-BART | 0.89 | 0.89 | 0.89 | ||
-CPT | 0.92 | 0.92 | 0.92 | ||
-Ours | 0.91 | 0.91 | 0.91 | ||
-4-class | dialog | BERT | 0.49 | 0.45 | 0.45 |
-BART | 0.53 | 0.53 | 0.52 | ||
-CPT | 0.49 | 0.47 | 0.46 | ||
-Ours | 0.57 | 0.56 | 0.53 | ||
-summary | BERT | 0.67 | 0.66 | 0.66 | |
-BART | 0.68 | 0.67 | 0.66 | ||
-CPT | 0.73 | 0.72 | 0.72 | ||
-Ours | 0.79 | 0.76 | 0.76 |
-Task | Input | Model | Precision | Recall | F1 |
-2-class | dialog | BERT | 0.81 | 0.78 | 0.79 |
-BART | 0.77 | 0.75 | 0.75 | ||
-CPT | 0.84 | 0.82 | 0.82 | ||
-Ours | 0.91 | 0.90 | 0.91 | ||
-4-class | dialog | BERT | 0.72 | 0.64 | 0.66 |
-BART | 0.70 | 0.66 | 0.65 | ||
-CPT | 0.76 | 0.68 | 0.70 | ||
-Ours | 0.79 | 0.71 | 0.73 |
+First, all models show substantial improvements in BLEU-2, ROUGE-L, and METEOR after applying POST, indicating that this method can significantly enhance the generation quality.
+Second, the application of POST resulted in considerable performance boosts across all models, with CPT showing the most significant improvement of +14.36% in BLEU-2 over its baseline.
+This enhancement is attributed to the framework’s ability to effectively assess the diagnostic stage, summarize patient information, and infer potential diagnoses.
+Further improvements are observed when golden POSTs are given.
+Third, as ChatGPT cannot be fine-tuned for our specific task, its baseline performance in this domain is restricted.
+However, the introduction of POST notably improved ChatGPT’s performance by +7.51% in BLEU-2, with substantial improvements upon using golden POST (+19.45%), demonstrating the effectiveness of the POST strategy.
+Interestingly, while ChatGPT shows improved DIST-2 with POST (+0.04), both CPT and ChatGLM3 exhibit slight decreases in DIST-2 (-0.01 and -0.06 respectively).
+This divergence reflects that unguided baselines naturally produce more diverse outputs, while POST’s structure constrains variety.
+ChatGPT’s unique characteristics allow it to maintain diversity even with structured guidance.
 5.3 Human Evaluation
@@
 Then, we assigned annotators with dialogue pairs to evaluate the performance of the doctor model following four aspects: 1) Fluency (Flu.) assesses the smoothness of the whole conversation; 2) Comforting (Com.) measures the ability to empathize and comfort; 3) Doctor-likeness (Doc.) gauges the adaptability in shifting topics based on the patient’s situation; 4) Engagingness (Eng.) measures if the model sustains attention throughout the conversation.
-As shown in Table 6, our model surpasses the baseline in the majority of aspects, particularly in doctor-likeness and engagingness, indicating significant improvements compared to other models.
-This suggests that incorporating the POST enhances our model’s ability to emulate doctors’ interactions, capturing the patient’s state in real-time, and allowing for flexible topic transitions throughout the dialogue.
-However, we also note that our model did not outperform ChatGPT in terms of comforting.
-This could be attributed to ChatGPT’s ability to generate extensive comforting and advisory responses without specific tuning.
-The dialogue examples of these models can be found in Appendix B.
-Comparisons | Aspect | Win | Lose | Tie |
-Ours vs. CPT | Flu. | 44.7 | 21.2 | 34.1 |
-Com. | 68.9 | 15.2 | 15.9 | |
-Doc. | 62.1 | 14.4 | 23.5 | |
-Eng. | 56.8 | 11.4 | 31.8 | |
-Ours vs. ChatGPT | Flu. | 39.4 | 34.8 | 25.8 |
-Com. | 33.3 | 55.3 | 11.4 | |
-Doc. | 53.8 | 34.1 | 12.1 | |
-Eng. | 41.7 | 28.8 | 29.5 |
+[…]
+ChatGLM3 +POST vs. ChatGLM3 | Flu. | 63.4 | 32.5 | 4.1 |
+Com. | 95.9 | 4.1 | 0.0 | |
+Doc. | 92.7 | 6.5 | 0.8 | |
+Eng. | 93.5 | 5.7 | 0.8 | |
+[…]
+As shown in Table 3, the baseline using the POST method performs better in most aspects, with particularly significant improvements in Comforting, Doctor-likeness, and Engagingness.
+Regarding fluency, the gap between our two models is not as large as other metrics, since the method has a relatively minor impact on fluency.
+These results suggest that incorporating the POST enhances our model’s ability to emulate doctors’ interactions, capturing the patient’s state in real-time, and allowing for flexible strategy transitions throughout the dialogue.
 5.4 Ablation Studies
@@
 The obtained results are demonstrated in Table 4.
-Model | BLEU-2 | ROUGE-L | METEOR | DIST-2 |
-Ours* | 45.28% | 0.56 | 0.5794 | 0.10 |
+The study reveals that omitting the next strategy (w/o Next) significantly impacts response generation, as it is directly correlated with the subsequent response’s relevance and effectiveness.
+Furthermore, excluding the stage (w/o Stage) notably affects the quality of generation, indicating the importance of the diagnostic dialogue phase in tailoring responses to be stage-appropriate and rational.
+In contrast, the absence of information (w/o Info) and summary (w/o Sum) components showed less impact on response generation quality, due to their roles in summarizing dialogue history rather than directly influencing responses.
+Nonetheless, these elements are crucial for their interpretability and utility in clinical settings, aiding doctors in summarizing patient symptoms and providing preliminary diagnostic insights, thereby facilitating further diagnosis and having a substantial role in clinical psychological consultations.
+[…]
+ChatGLM3 +POST* | 45.28% | 0.56 | 0.5794 | 0.10 |
 w/o Info | 44.43% | 0.54 | 0.5708 | 0.16 |
@@
 w/o Next | 36.24% | 0.45 | 0.5107 | 0.18 |
-The study reveals that omitting the next topic (w/o Next) significantly impacts response generation, as it is directly correlated with the subsequent response’s relevance and effectiveness.
-This omission led to an increase in the DIST-2 value, suggesting a reduction in response content diversity due to the focused nature of topic-specific responses.
-Furthermore, excluding the stage (w/o Stage) notably affects the quality of generation, indicating the importance of the diagnostic dialogue phase in tailoring responses to be stage-appropriate and rational.
-In contrast, the absence of information (w/o Info) and summary (w/o Sum) components showed lesser impact on response generation quality, due to their roles in summarizing dialogue history rather than directly influencing responses.
-Nonetheless, these elements are crucial for their interpretability and utility in clinical settings, aiding doctors in summarizing patient symptoms and providing preliminary diagnostic insights, thereby facilitating further diagnosis and having a substantial role in clinical psychological consultations.
+[…]
 6 Conclusions
 In this paper, we incorporate Psychological State Tracking (POST) within LLM to guide response generation for doctors during depression diagnosis consultations.
-In particular, the state is defined based on the Albert Ellis’ ABC Model in psychology, which illuminates a profound connection between patients’ information changes and doctors’ strategic planning.
+In particular, the state is defined based on Albert Ellis’ ABC Model in psychology, which illuminates a profound connection between patients’ information changes and doctors’ strategic planning.
 Extensive experiments show that the integration of psychological state tracking significantly enhances the performance of LLMs to generate responses in depression-diagnosis-oriented chat.
@@
 In the future, we will explore more specialized and fine-grained state tracking methods and incorporate patient personalized information to guide diagnosis-oriented chat.
-Limitations
-The main limitation of our approach lies in the evaluation metrics.
-In the field of diagnostic chat, traditional automatic evaluation methods may not accurately reflect the true effectiveness of the models.
-As highlighted by Liu et al.
-(2016), there exists a certain discrepancy between the results and human evaluations.
-Developing more comprehensive, considerate, and task-specific evaluation metrics remains a challenge to be addressed in future research.
-Ethics Statement
-In this study, we annotate the psychological state collaboratively by professional labors and GPT-4, followed by a thorough data cleaning and filtering process.
-This step ensured the removal of any private or sensitive information from the dataset.
-Acknowledgements
-We acknowledge the contributions of the SJTU X-LANCE Lab555https://x-lance.sjtu.edu.cn/en in constructing the original depression-diagnosis-oriented chat dataset (i.e., D).
-We cite the work and conduct further research by adapting the dataset and exploring LLM-based method from the psychological views.
-Appendix A Prompt Template
-Description: I want you to serve as a psychological counselor, and use rational emotive behavior therapy and ABC theory to mark the current patient situation and the next strategy.
-Please mark the input dialogues according to the requirements and examples in Chinese.
-The sequence numbers of the input dialogues must strictly correspond one-to-one with those of the outputs, with no merging or omissions allowed.
-Requirements: Definitions and marking requirements of four parts of POST: <Stage> The current diagnostic stage (Activating Event (A) / Belief (B) / Consequence (C), please enter the complete stage, and answers like "same as above" are not allowed).
-<Info> Current patient information (keep information from historical conversations as briefly as possible).
-<Sum> Diagnosis result of the patient’s condition (you need to refine it to the underlying disease).
-<Next> Next step strategy (choose the most suitable option from 11 strategies).
-Examples: Examples pre-labeled by humans Input: Input data Output:
-Appendix B Case Study
-Figures 5, 6, 7 respectively display dialogue examples of our model, CPT, and ChatGPT under the same patient background, where these models play the role of doctors.
-In the diagram, statements of empathetic consolation are highlighted with green wavy lines, while inquiries of diagnosis are underlined in blue.
-It is evident from the illustration that, in comparison to other models, our model offers more natural empathetic consolation and topics of inquiry that are closely aligned with the diagnostic state.
-The responses from CPT tend to be relatively rigid, with less empathetic consolation and more straightforward, singular diagnostic inquiries.
-Although ChatGPT provides commendable empathetic consolation, its replies often are too lengthy, containing extensive analysis and advice, yet they do not closely correlate with the diagnostic content, lacking the distinctive characteristics of a doctor.
+6.0.1 Acknowledgement
+This research is funded by the National Nature Science Foundation of China (No.
+62477010 and No.62307028), the Natural Science Foundation of Shanghai (No.
+23ZR1441800), Shanghai Science and Technology Innovation Action Plan (No.
+24YF2710100 and No.23YF1426100) and Shanghai Special Project to Promote High-quality Industrial Development (No.
+2024-GZL-RGZN-02008).